### v1

# Knowledge Graph Pruning for Recommendation

Abstract.
Recent years have witnessed the prosperity of knowledge graph based recommendation system (KGRS), which enriches the representation of users, items, and entities by structural knowledge with striking improvement. Nevertheless, its unaffordable computational cost still limits researchers from exploring more sophisticated models. We observe that the bottleneck for training efficiency arises from the knowledge graph, which is plagued by the well-known issue of knowledge explosion. Recently, some works have attempted to slim the inflated KG via summarization techniques. However, these summarized nodes may ignore the collaborative signals and deviate from the facts that nodes in knowledge graph represent symbolic abstractions of entities from the real-world. To this end, in this paper, we propose a novel approach called KGTrimmer for knowledge graph pruning tailored for recommendation, to remove the unessential nodes while minimizing performance degradation. Specifically, we design an importance evaluator from a dual-view perspective. For the collective view, we embrace the idea of collective intelligence by extracting community consensus based on abundant collaborative signals, i.e. nodes are considered important if they attract attention of numerous users. For the holistic view, we learn a global mask to identify the valueless nodes from their inherent properties or overall popularity. With the collective and holistic importance scores, we build an end-to-end importance-aware graph neural network, which injects filtered knowledge to enhance the distillation of valuable user-item collaborative signals. Ultimately, we generate a pruned knowledge graph with lightweight, stable, and robust properties to facilitate the following-up recommendation task. Extensive experiments are conducted on three publicly available datasets to prove the effectiveness and generalization ability of KGTrimmer, where it can reduce the number of triplets in KG by up to 90% without compromising performance.
Knowledge Graph Pruning, Recommender Systems, Graph Neural Networks
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. INTRODUCTION
Incorporating knowledge graph (KG) as a supplementary resource to support the recommendation system has gained significant popularity. By leveraging KGs, we can obtain a more comprehensive understanding among users, items, and entities, which leads to a better characterization of user preference and item relevance (Wang et al., 2021a, 2019a; Du et al., 2022). KGs can also enhance the transparency and interpretability of recommendation systems, as they provide clear evidence of how we arrive at the recommendation results (Wang et al., 2021a).
To better leverage the KG data, numerous researchers are dedicated to incorporating users, items, and entities into a shared embedding space based on a collaborative knowledge graph (CKG), which aims to manipulate the structured knowledge and produce explainable recommendation results (Wang et al., 2019a). This line of literature adopts the widespread message aggregation scheme by recursively collecting neighborhood information on the CKG to capture the long-range connectivities (Zhou et al., 2020; Wang et al., 2019a; Yang et al., 2023; Wang et al., 2021a). Hence, the valuable knowledge information could be distilled from the KG to boost the collaborative signals over the user-item interactions.
Figure 1. The motivation example from an online a movie platform, where (a) presents a real-world collaborative knowledge graph and (b) indicates the corresponding collective importance of each entity for user modeling.
Despite the prevalence, previous literature is based on the steadfast assumption that KGs consistently deliver accurate, reliable, and valuable facts, thereby undoubtedly enhancing recommendation performance (Wang et al., 2021a; Du et al., 2022; Yang et al., 2023). However, in real-world scenarios, large-scale hand-crafted KGs frequently suffer from evident knowledge redundancy and encounter uncontrolled knowledge quality (Xu et al., 2020; Wang et al., 2021b). This not only imposes additional computational overhead but may even adversely undermine recommendation performance. Inspired by this, pruning the KGs is indispensable to better serve recommendation systems with improvements on both efficiency and accuracy.
To this end, we investigate the knowledge graph pruning approach tailored for recommendation systems, which is a never-before-studied problem. Early efforts has been devoted to KG summarization (Doerr and Blenn, 2013; Sacenti et al., 2022; Kang et al., 2022; Brody et al., 2022; Goyal et al., 2020), employing techniques such as cluster-based inference (Doerr and Blenn, 2013; Sacenti et al., 2022) or GNN-based approaches (Kang et al., 2022; Brody et al., 2022; Goyal et al., 2020) to compress the KGs and enhance subsequent utilization efficiency. However, this line of research shares inherent limitations, as the summarized nodes may deviate from the facts that nodes in knowledge graph represent symbolic abstractions of entities from the real-world, thereby impacting the interpretability of the KG. On the contrary, KG pruning aims to obtain a learnable vector to determine the preservation or removal of each triplet, so that we can evaluate their contribution while preserving the original semantics of knowledge. Furthermore, studies in this field solely evaluate the importance of individual attributes or triplets from the structure of KG, without any consideration of recommendation task. Consequently, the pruning result cannot precisely align with the recommendation system, resulting in suboptimal performance.
In this paper, we target at eliminating the valueless nodes in the KG with respect to recommender system. Formally, we define the nodes that cannot bring any information gain for KG-based recommendation task as valueless nodes. In order to tackle this task, there are two challenges urged to be addressed:
  • How to evaluate the importance of nodes? Traditional pruning methods emphasize the structural coherence and connectivity within knowledge graphs, yet such methods are inadequately suited for pruning tasks in KG-based recommendation systems. For example, in conventional pruning tasks, nodes with high degrees are more likely to be preserved , as their removal would greatly compromise the connectivity of the KG (Faralli et al., 2018). In contrast, when it comes to KG-based recommendation tasks, things may be different: collaborative filtering signals are introduced but the majority of literature (Faralli et al., 2018; Doerr and Blenn, 2013; Zhang et al., 2022) fails to integrate it to identify the knowledge that is useless for the accurate modeling of user profiles. Taking Fig 1(a) as an example, as all movies are connected to ”English” node, it is incapable of differentiating user preferences and therefore contributes less informational value to the recommender tasks. Hence, we need to formulate an innovative importance criterion that incorporates insights from collaborative filtering signals.
  • How can we acquire ground truth to inform and direct the pruning process? To the best of our knowledge, the mainstream of the works shows limited adaptability to the task. Supervised methods primarily rely on manually introduced erroneous nodes as labels (Zhang et al., 2022). However, although randomly injected nodes could be perceived as noises or errors, they can hardly be regarded as valueless nodes. Moreover, Unsupervised studies(Belth et al., 2020) utilize rule-learning approaches to detect strange or abnormal facts. Nevertheless, applying these method is ill-advised, as the valueless nodes could be retained due to their high occurrence. Fortunately, Some recent studies (Xu et al., 2022) attempt to utilize the gold answers from the downstream tasks to guide the pruning process. In light of this, we are expected to devise a strategy to deliver the signals to pruning tasks.
To face the challenges, we propose an adaptive knowledge graph pruning model tailored for recommendation called KGTrimmer, aiming to remove uninformative or valueless knowledge while minimizing performance degradation. Initially, we evaluate the contribution of each entity in the KG through an importance evaluator from a dual-view perspective. For the collective view, we embrace the idea of collective intelligence by extracting and aggregating consensus based on collaborative signals, i.e. nodes are considered important if they attract attention of numerous users. As depicted in 1 (b), neither ”James” nor ”Alice” shows interest on ”English”, then ”English” will be tagged as valueless node. For the holistic view, we obtain the holistic importance of each entity via a global mask to identify the valueless nodes from their inherent properties or overall popularity. By aggregating the importance scores of the collective and holistic views, we establish an aggregated mask vector that automatically identifies the essential and indispensable entities in the KG. Afterward, we build an end-to-end importance-aware graph neural network, which simultaneously exploits user-item interactions and the item-item relevance with the filtering mechanism on the CKG. Upon identifying the important nodes with the iterative process, we ultimately generate a valuable, stable, and robust KG after careful knowledge pruning in the context of the given recommendation scenarios. We make the following main contributions:
  • We are among the first ones to develop knowledge graph pruning approaches specifically tailored for personalized recommendation scenarios, intending to expedite the training process while not compromising predictive accuracy.
  • We propose a novel model called KGTrimmer. Specifically, we automatically select the essential and valuable entities in a dual-view perspective, i.e., collective and holistic views. Then, we design an importance-aware graph neural network with constraint information flow to distill the informative knowledge for user modeling.
  • We conduct extensive experiments on three public benchmarking datasets to validate the superiority of our proposed KGTrimmer. The pruned KG preserves the most essential facts and better serves the recommendation system.
2. PRELIMINARIES
In this section, we will begin by introducing the concepts of knowledge graph and KG-based recommendation systems. Then, we will proceed to provide a formal definition of knowledge graph pruning specifically for recommendation systems.
2.1. KG-based Recommendation System
A knowledge graph serves as a repository that stores structured information about real-world facts in the form of a heterogeneous graph, encompassing item attributes, taxonomy, and external commonsense knowledge.
Definition 2.1.
Knowledge Graph (KG). Given an entity set $\mathcal{V}$ and a relation set $\mathcal{R}$, a knowledge graph can be denoted as $G=\{(h,r,t)\ \mid h,t\in\mathcal{V},r\in\mathcal{R}\}$. In particular, each triplet $(h,r,t)$ in the KG represents a connection between the head entity $h$ and the tail entity $t$, which is linked by a specific relation $r$. It is noteworthy that $\mathcal{R}$ is comprised of relations from both canonical and inverse directions.
In recommendation scenarios, KGs have been proven to significantly enhance the intricate relevance among items and extract deep-seated user preferences, thereby improving recommendation performance. Following traditional works, suppose there exists a user set $\mathcal{U}$ ($|\mathcal{U}|=M$) and an item set $\mathcal{I}$ ($|\mathcal{I}|=N$), the historical interactions between users and items (e.g., click or purchase) can be naturally transformed into a user-item bipartite graph. Specifically, the binary interaction matrix is denoted as $Y\in\{0,1\}^{M\times N}$, where each entry $y_{ui}=1$ indicates that the given user $u\in\mathcal{U}$ has interacted with item $i\in\mathcal{I}$, otherwise $y_{ui}=0$. Furthermore, we establish an item-entity collection for alignment, denoted as $\mathcal{A}=\{(i,e)\mid i\in\mathcal{I},e\in\mathcal{V}\}$, which allows us to align the item $i$ with the entity $e$. Hence, we seamlessly integrate the KG with the user-item bipartite graph and subsequently develop a unified collaborative knowledge graph (CKG) for the downstream tasks.
Definition 2.2.
KG-based Recommendation System (KGRS). Given the user set $\mathcal{U}$, the item set $\mathcal{I}$, and their observed interactions $Y$, with the complementary knowledge from $G$, the target of the KGRS task is to learn a function $\mathcal{F}$ capable of predicting the probability that a user $u\in\mathcal{U}$ interacts with a specific item $i\in\mathcal{I}$.
2.2. Knowledge Graph Pruning for Recommendation
Here we present the definition of knowledge graph pruning for recommendation task (KGPR) which targets at reducing redundant and uninformative edges in knowledge graph $G$ for efficient and effective prediction in KGRS. Specifically, the objective of KGPR task is to learn a binary mask vector $\boldsymbol{s}\in\{0,1\}^{|G|}$ for all triplets, where the $k$-th edge $(h_{k},r_{k},t_{k})$ will be preserved if ${s}_{k}=1$, otherwise removed. To this end, we employ the binary vector $\boldsymbol{s}$ on the original KG to obtain the pruned KG, denoted as $\tilde{G}=G\odot\boldsymbol{s}$. To be specific, the formal definition of KGPR task can be described as:
Definition 2.3.
Knowledge Graph Pruning for Recommendation Systems (KGPR). Given a user set $\mathcal{U}$, an item set $\mathcal{I}$, an interaction matrix $Y$ and a knowledge graph $G$, we aim to learn a binary function $\mathcal{H}$, s.t. $\boldsymbol{s}=\mathcal{H}(\mathcal{U},\mathcal{I},Y,G)$ and $\boldsymbol{s}\in\{0,1\}^{|G|}$, where $\boldsymbol{s}$ is a binary mask vector. Subsequently, we can acquire the pruned KG, denoted as $\tilde{G}=G\odot\boldsymbol{s}$. According to the pruned KG, the objective of KGPR is to maximize the predictive accuracy of the unobserved user-item interactions in KGRS.
Figure 2. The overall framework of the proposed KGTrimmer framework, which consists three components: (a)Dual-View Importance Evaluator, (b) Importance-Aware Graph Neural Network, and (c) Pruned Knowledge Graph Generation.
3. METHODOLOGY
In this section, we will explain the KGTrimmer model to achieve knowledge graph pruning for recommendation in detail, which is composed of the following modules, as shown in Figure 2:
  • Dual-view Importance Evaluator. We assess the importance of the candidate entities from a dual-view perspective: collective and holistic view. On one hand, the collective view explicitly utilizes the user attention towards entities to evaluate the importance of entities. On the other hand, the holistic view focuses on the role of the entity within the overall recommendation task. Both contribute to the identification of valueless nodes.
  • Importance-Aware Graph Neural Network. We leverage the ultimate importance scores to control the information flow for general neural network propagation. In this way, this refined approach allows for the transmission of signals from the recommendation system to the dual-view evaluator, where the module can prioritize crucial information and discard the redundant one.
  • Pruned Knowledge Graph Generation. We finally aggregate the candidate masks during the training process to obtain a stable and robust pruned knowledge graph.
3.1. Dual-view Importance Evaluator
In this subsection, we will introduce how we construct the dual-view KGTrimmer to filter out the effective knowledge that is really beneficial to the collaborative filtering. Specifically, we derive the importance score for each entity from two views, i.e. collective view and holistic view. For the collective view, we try to consider the user-item interactions to compute the collective importance scores, which show the specific interacted users’ preference on the entities. For the holistic view, a learnable mask vector is applied which expects to identify the valueless entities from inherent properties or overall popularity, which is known as the holistic importance scores. After that, we balance the influence of the collective and holistic views and aggregate their importance scores to obtain the aggregated importance scores on all KG entities.
3.1.1. Collective View
Users express personalized preferences over items while the items are linked with the entities. Along this line, we realize that the essence of user interest lies in the entities within the KG. Motivated by the concept of collective intelligence, we take advantage of these abundant user-item interactions and consider them as the results of a social discourse. Coupled with the KG, we collect and aggregate their diverse opinions to extract the consensus and determine the importance of each entity for shaping their user profiles in a collective view. To this end, we intuitively construct a real-value based user-entity matrix to better represent the correlation between the users and entities. After the construction, each entity receives the attention from all related users to generate a score, a.k.a. collective score. We delve into the details as follows:
User-Entity Matrix Construction. For a more streamlined computational process, we gather data from collaborative signals and knowledge graph to assemble the user-entity matrix. To be more specific, we first follow the prior studies (Wang et al., 2019a) to build a collaborative knowledge graph by combining the user-item bipartite graph and the knowledge graph. For simplicity, we denote the matrix as $Q\in\{0,1\}^{|\mathcal{U}|\times|\mathcal{V}|}$, where $|\mathcal{U}|$ and $|\mathcal{V}|$ is the number of users and entities respectively. Then we define :
(1) $Q_{uv}=\begin{cases}1,&\text{ there is a path from user }u\text{to entity }v\\ 0,&\text{ otherwise }\end{cases}$
Due to the knowledge explosion, an entity may reach uncountable users, thereby leading to the substantial computational cost. As a result, we employ a more effective strategy to consider the path from user $u$ to entity $v$. That is to say, if an entity $v$ maintains more than $k$ users, i.e. $|Q_{v}|>k$, we implement a random sampling strategy to set $Q_{uv}=1$ for the $k$ selected users, while setting $Q_{uv}=0$ for the rest. Here $|Q_{v}|$ represents the number of paths that end with entity $v$.
Collective Importance Scores. In this part, we target at revealing the collective importance of the entities in the KG through user-entity matrix. As mentioned above, for each entity $v$, we first refer to the $Q$ matrix and denote the user neighbors of the entity $v$ as $\mathcal{N}_{v}^{Q}=\{u|Q_{uv}=1,u\in\mathcal{U}\}$. Afterward, we collect the opinions of interacted users to compute the collective importance score for the specific entity. Besides, we also enrich the representation of entities by incorporating the information of relation. The formulation can be described as follows:
(2) $s_{v}^{c}$ $=\frac{1}{|\mathcal{N}_{v}^{Q}|}\sum_{u\in\mathcal{N}_{v}^{Q}}sim(\boldsymbol{e}_{u},\boldsymbol{e}_{v}^{\prime}),$
$\boldsymbol{e}_{v}^{\prime}$ $=\frac{1}{|\mathcal{N}_{v}^{\mathcal{R}}|}\sum_{r\in\mathcal{N}_{v}^{\mathcal{R}}}\boldsymbol{e}_{r}\odot\boldsymbol{e}_{v},$
where $\mathcal{N}_{v}^{\mathcal{R}}=\{r|(h,r,v)\in G\}$, $sim$ is the similarity function, $\odot$ denotes the element-wise product, $\boldsymbol{e}_{u}$, $\boldsymbol{e}_{v},\boldsymbol{e}_{r}$ are the embeddings of user $u$, entity $v$ and relation $r$, respectively, and $\boldsymbol{e}_{v}^{\prime}$ is the entity embedding which absorbs the information of relations, and $s_{v}^{c}$ is the collective importance score for each entity $v$. Formally, we define the similarity function as :
(3) $sim(\boldsymbol{e}_{u},\boldsymbol{e}_{v}^{\prime})$ $=max(0,\frac{\boldsymbol{e}_{u}^{T}\boldsymbol{e}_{v}^{\prime}}{\|\boldsymbol{e}_{u}\|\|\boldsymbol{e}_{v}^{\prime}\|}),$
We use the modified cosine similarity function for two reasons: (1) We can constrain the output to be within the range of 0 to 1, allowing it to be utilized as a score. (2) Entities with negative similarity are deemed noisy or erroneous nodes, therefore will be assigned the score of zero and prioritized for removal. Therefore, the collective score is endowed with the ability to jointly identify both noisy and valueless nodes. Consequently, we obtain the collective importance scores as:
(4) $\boldsymbol{s}^{c}=(s_{1}^{c},\dots,s_{|\mathcal{V}|}^{c}),$
which significantly demonstrates interacted users’ attention on KG entities based on collaborative signals.
3.1.2. Holistic View
The task of KG-based recommendation involves various aspects, such as purchase motivation, product popularity, degree of nodes, etc. Thus the KG-based recommender system inherently has the potential to identify the holistic importance of entities. To explicitly model this, we assign a learnable parameter to each entity, dubbed as holistic importance score.
Holistic Importance Scores. Similar to (Chen et al., 2021), we define a global mask to represent the underlying value of entities within the KG where the holistic importance scores can be formulated as:
(5) $\boldsymbol{s}^{h}=(s_{1}^{h},\dots,s_{|\mathcal{V}|}^{h}),$
Notably, we still limit the range of the holistic importance score between 0 and 1. By leveraging the advantages of the recommender system, the holistic scores possess the capacity to represent the significance of the entities. In other words, entities with low importance scores are ineffective at capturing valuable information for the recommendation task, whereas informative entities with high importance scores are the key drivers in enhancing the performance of the recommendation system.
3.1.3. Aggregated Importance Scores
To achieve the knowledge pruning for recommendation, we are desired to perform aggregation operations on these two scores. Formally, the aggregated importance scores for entities can be represented as:
(6) $\hat{\boldsymbol{s}}=\gamma\boldsymbol{s}^{c}+(1-\gamma)\boldsymbol{s}^{h}$
where $\gamma$ is a trade-off hyper-parameter. In this way, $\boldsymbol{\hat{s}}$ can be applied for all entities so as to discover the unnecessary or uninformative entities that are expected to be pruned from the KG. Subsequently, we design a transformation function $\mathcal{T}$ to to measure each triplet $(h,r,t)$ according to the acquired $\boldsymbol{\hat{s}}$:
(7) $\boldsymbol{\hat{s}}(h,r,t)=\mathcal{T}(\boldsymbol{\hat{s}}_{h},\boldsymbol{\hat{s}}_{t}).$
where the specific form will be discussed in the Experimental Section.
3.2. Importance-Aware Graph Neural Network
In this part, we aim to establish the connection between the recommendation system and the evaluator module. In detail, we inject the final importance scores into the message propagation process of formal KGRS models and make predictions on the unobserved user-item interactions. Previous Graph-based approaches have shown that the neighborhood-based aggregation scheme is a promising method to capture the user preference with the help of KG, such as KGIN (Wang et al., 2021a) and KGRec (Yang et al., 2023). Specifically, after broadcasting the importance scores of entities to the corresponding triplets, we emphasize that the key point turns to incorporating the aggregated importance scores into the message propagation process through a dedicated graph neural network and learning the refined representations of users, items, and entities on the CKG. To be specific, we illustrate knowledge graph propagation and user-item graph propagation in detail.
3.2.1. Knowledge Graph Propagation
We move on to learning the representations of items and entities with the aggregated importance scores $\boldsymbol{\hat{s}}$. Since some valueless or noisy entities hurt the accuracy or efficiency of KGRS, the importance scores initially provide guidance for knowledge filtering. Formally, we iteratively receive and update the embeddings with constrained information flow in terms of the importance scores as follows:
(8) $\boldsymbol{e}_{v}^{(l)}=KGE(\boldsymbol{e}_{v}^{(l-1)},\boldsymbol{e}_{r},\boldsymbol{\hat{s}}).$
Here, we denote $KGE$ as any KG embedding method for entity embedding update, $\boldsymbol{e}_{v}^{(l-1)}$ and $\boldsymbol{e}_{v}^{(l)}$ are entity embeddings of previous layer $l-1$ and current layer $l$, respectively. In detail, we adopt the following approach for knowledge graph propagation:
(9) $\boldsymbol{e}_{v}^{(l)}=\frac{1}{|\mathcal{N}_{v}^{G}|}\sum_{(r,t)\in\mathcal{N}_{v}^{G}}\boldsymbol{\hat{s}}(v,r,t)(\boldsymbol{e}_{t}^{(l-1)}\odot\boldsymbol{e}_{r}).$
Here $\mathcal{N}_{v}^{G}=\{(r,t)|(v,r,t)\in G\}$ includes the tail entities with their corresponding relation set for a given entity $v$.
3.2.2. User-Item Graph Propagation
With the integration of KG, we can delve more profoundly into users’ preferences for entities in the KG, which provides more convincing evidence for the recommendation system. Hence, we attempt to apply the filtered knowledge information to enhance the distillation of valuable user-item collaborative signals, leading to a more accurate understanding of user preferences. With the refined item embeddings by knowledge graph propagation, the user embeddings can be calculated as:
(10) $\boldsymbol{e}_{u}^{(l)}=CF(\boldsymbol{e}_{u}^{(l-1)},\boldsymbol{e}_{i}^{(l-1)}).$
Here $CF$ can be any collaborative filtering method to update user embeddings. $\boldsymbol{e}_{u}^{(l-1)}$ and $\boldsymbol{e}_{u}^{(l)}$ are the user embeddings of previous layer $l-1$ and current layer $l$, and $\boldsymbol{e}_{i}^{(l-1)}$ is the item embedding of previous layer $l-1$. Specifically, we simply adopt the following aggregation method to update user embeddings:
(11) $\boldsymbol{e}_{u}^{(l)}=\frac{1}{|\mathcal{N}_{u}|}\sum_{i\in\mathcal{N}_{u}}\boldsymbol{e}_{i}^{(l-1)},$
where $\mathcal{N}_{u}=\{i|Y_{ui}=1\}$ are the item neighbors of user $u$.
3.2.3. Prediction
After stacking $L$ layers, the user embedding $\boldsymbol{e}_{u}$ and item embedding $\boldsymbol{e}_{i}$ can be aggregated as:
(12) $\boldsymbol{e}_{u}$ $=\boldsymbol{e}_{u}^{(0)}+\boldsymbol{e}_{u}^{(1)}+\dots+\boldsymbol{e}_{u}^{(L)}$
$\boldsymbol{e}_{i}$ $=\boldsymbol{e}_{i}^{(0)}+\boldsymbol{e}_{i}^{(1)}+\dots+\boldsymbol{e}_{i}^{(L)}.$
where $L$ denotes the number of layers of message propagation. Then we adopt inner product to compute predictive scores and utilize BPR loss to train the recommendation model:
(13) $\mathcal{L}=\sum_{u\in\mathcal{U}}\sum_{i\in\mathcal{N}_{u}}\sum_{j\notin\mathcal{N}_{u}}^{N_{neg}}\log\left(\sigma\left(\boldsymbol{e}_{u}^{\top}\boldsymbol{e}_{i}-\boldsymbol{e}_{u}^{\top}\boldsymbol{e}_{j}\right)\right),$
where $N_{neg}$ is the number of negative samples for each user-item interaction in the training set.
3.3. Pruned Knowledge Graph Generation
In this section, a final pruned knowledge graph $\tilde{G}$ will be generated for the downstream tasks based on $\boldsymbol{\hat{s}}(h,r,t)$. In fact, during the training process, the candidate mask $\boldsymbol{\hat{s}}^{k}$ can be generated in each epoch. At last, the objective mask vector $\boldsymbol{\tilde{s}}$ can be presented as:
(14) $\boldsymbol{\tilde{s}}=AGG(\boldsymbol{\hat{s}}^{0},\boldsymbol{\hat{s}}^{1}\dots,\boldsymbol{\hat{s}}^{K}),$
where $K$ is number of training epochs, $AGG$ is the function to aggregate all learned mask vectors of each epoch. Additionally, to map the scores to $\{0,1\}$, we take $\boldsymbol{s}=\boldsymbol{1}(\boldsymbol{\tilde{s}}\geq\tau)$ to mask edges in knowledge graph $G$, here $\boldsymbol{1}$ is the indicator function and $\tau$ is a threshold. Consequently, the pruned KG can be constructed as:
(15) $\tilde{G}=G\odot\boldsymbol{s}$
3.4. Model Complexity Analysis
The time cost mainly comes from two parts: collective importance score calculation and Importance-Aware GNN module. In the first part, the time expense is greatly reduced because we randomly sampled $k$ users before the calculation. Thus the computational complexity is $O(k|\mathcal{V}|d)$, where $|\mathcal{V}|$ is the number of entities and $d$ is the embedding size. The second part can further be divided into two phase: collaborative filter phase and KG aggregation phase. As for collaborative filter phase, the time cost is $O(L|Y|d)$, where $L$ is the number of layer and $|Y|$ is the number of interaction, respectively. Moreover, the computational consumption of KG aggregation phase is $O(L\mathcal{|G|}d)$, where $|\mathcal{G}|$ is the number of triplets. Notably, compared with the cost of the Importance-Aware GNN module, the expense of collective importance score calculation is negligible.
4. EXPERIMENTS
4.1. Experimental Settings
4.1.1. Data Description.
To demonstrate the superiority of the proposed KGTrimmer framework, we conduct experiments on three publicly available benchmarking datasets for KGRS, namely Last-FM and Alibaba-Fashion, for comprehensive investigation. The statistics of datasets are summarized in Table 1.
  • Amazon-book [^1] http://jmcauley.ucsd.edu/data/amazon.: This is a popular amazon dataset for product recommendation. In addition, we refer to (Zhao et al., 2018) and construct the KG data by mapping items into Freebase entities.
  • Last-FM [^2] https://grouplens.org/datasets/hetrec-2011.: This is a widely-adopted music record dataset released by Last.fm platform, where the tracks are regarded as items. Particularly, the data ranges from Jan 2015 to June 2015 is collected for experiments. Similarly, we align the items with the Freebase entities to fetch the KG data.
  • Alibaba-Fashion [^3] https://github.com/huangtinglin/Knowledge_Graph_based_Intent_Network.: This is a fashion outfit dataset collected from Alibaba which is an online shopping platform. The outfits are recommended to users as items, where each outfit consists of several fashion stuffs(e.g., tops, bottoms, shoes), and these stuffs are assigned different fashion categories(e.g., jeans, T-shirts). And these attributes are constructed as the KG data of outfits.
To ensure the data quality, we use the 10-core setting to guarantee that each user or item has at least 10 interactions. Similarly, the infrequent entities (less than 10 triplets) and relations (less than 50 triplets) are filtered out for the KG part of both datasets. It is worth noting that we utilize the same processed datasets released by KGIN (Wang et al., 2021a) and KGAT (Wang et al., 2019a), respectively.
Table 1. Statistics of three real-world datasets.
Statistics Amazon-book Last-FM Alibaba-Fashion
User-Item Interactions # Users 70679 23,566 114,737
# Items 24915 48,123 30,040
# Interacts 847733 3,034,796 1,781,093
Knowledge Graph # Entities 88572 58,266 59,156
# Relations 39 9 51
# Triplets 2557746 464,567 279,155
4.1.2. Evaluation Metrics.
In the evaluation phase, we follow the previous studies and adopt the all-ranking strategy for convincing experiments. To measure the performance, two representative evaluation metrics for top-K recommendation, i.e., $Recall@K$ and $NDCG@K$ (Normalized Discounted Cumulative Gain) are adopted for all experiment settings. Note we set $K=20$ by default.
Table 2. The experimental comparison among a wide range of knowledge graph pruning methods for three benchmarking datasets. Note that KGIN and KGRec are selected KG-based recommendation models as the backbone. R@20 and N@20 stand for Recall@20 and NDCG@20, respectively. The best results are in bold and the secondary best results are underlined.
Dataset Backbone Ratio Metric Pop KGRS-IMP RotatE KGAT CAGED KGist KGTrimmer
Amazon-book KGIN 90% R@20 0.1234 0.1647 0.1573 0.1621 0.1132 0.1196 0.1667
N@20 0.0648 0.0890 0.0850 0.0871 0.0587 0.0636 0.0898
95% R@20 0.1285 0.1460 0.1560 0.1620 0.1204 0.1198 0.1662
N@20 0.0683 0.0775 0.0838 0.0871 0.0627 0.0637 0.0890
98% R@20 0.1270 0.1246 0.1423 0.1508 0.1260 0.1219 0.1615
N@20 0.0670 0.0648 0.0753 0.0803 0.0659 0.0636 0.0861
KGRec 90% R@20 0.1298 0.1549 0.1555 0.1561 0.1113 0.1141 0.1620
N@20 0.0685 0.0821 0.0837 0.0834 0.0573 0.0603 0.0861
95% R@20 0.1236 0.1443 0.1542 0.1563 0.1189 0.1138 0.1626
N@20 0.0650 0.0757 0.0821 0.0836 0.0618 0.0600 0.0860
98% R@20 0.1240 0.1225 0.1410 0.1460 0.1179 0.1120 0.1557
N@20 0.0647 0.0640 0.0745 0.0774 0.0612 0.0573 0.0826
Last-FM KGIN 30% R@20 0.0965 0.0964 0.0965 0.0922 0.0955 0.0932 0.0966
N@20 0.0843 0.0849 0.0835 0.0799 0.0825 0.0815 0.0838
70% R@20 0.0848 0.0857 0.0899 0.0830 0.0893 0.0866 0.0938
N@20 0.0738 0.0749 0.0782 0.0714 0.0773 0.0753 0.0824
90% R@20 0.0730 0.0791 0.0790 0.0738 0.0803 0.0798 0.0889
N@20 0.0626 0.0703 0.0704 0.0647 0.0689 0.0687 0.0789
KGRec 30% R@20 0.0927 0.0926 0.0922 0.0884 0.0906 0.0905 0.0941
N@20 0.0811 0.0809 0.0799 0.0764 0.0778 0.0785 0.0809
70% R@20 0.0838 0.0788 0.0863 0.0790 0.0837 0.0843 0.0909
N@20 0.0718 0.0695 0.0742 0.0681 0.0722 0.0728 0.0793
90% R@20 0.0775 0.0798 0.0753 0.0780 0.0805 0.0781 0.0870
N@20 0.0654 0.0690 0.0655 0.0674 0.0688 0.0669 0.0762
Alibaba-Fashion KGIN 30% R@20 0.1163 0.1171 0.1143 0.1070 0.1107 0.1169 0.1179
N@20 0.0724 0.0731 0.0712 0.0648 0.0686 0.0729 0.0738
70% R@20 0.1112 0.1062 0.1098 0.1018 0.0967 0.1051 0.1158
N@20 0.0690 0.0643 0.0679 0.0615 0.0588 0.0644 0.0725
90% R@20 0.1028 0.0975 0.1025 0.0970 0.0936 0.0965 0.1112
N@20 0.0627 0.0590 0.0627 0.0583 0.0562 0.0579 0.0694
KGRec 30% R@20 0.1194 0.1199 0.1176 0.1184 0.1141 0.1185 0.1199
N@20 0.0749 0.0749 0.0734 0.0741 0.0707 0.0741 0.0752
70% R@20 0.1114 0.1072 0.1101 0.1016 0.1000 0.1072 0.1165
N@20 0.0692 0.0654 0.0679 0.0614 0.0611 0.0658 0.0729
90% R@20 0.1027 0.0994 0.1010 0.0978 0.0925 0.0985 0.1104
N@20 0.0625 0.0605 0.0611 0.0590 0.0552 0.0589 0.0690
4.1.3. Backbones.
To demonstrate the effectiveness and efficiency of the pruned knowledge graph, we employ the following widely-adopted knowledge graph-based recommendation models as the backbone.
  • KGIN (Wang et al., 2021a): This work considers user-item interactions at the finer granularity of intents that are a combination of KG relations for better model capacity and interpretability for the recommendation. To this end, the proposed KGIN utilizes a relational path-aware aggregation module to identify the influences of different intents.
  • KGRec (Yang et al., 2023): It is a generative and contrastive self-supervised method to enhance recommender systems by KG. Specifically, KGRec proposes a rational-aware masking mechanism to identify the most important information in the KG, which suppresses potential noise and irrelevant knowledge graph connections.
4.1.4. Compared baselines.
Moreover, several pruning methods are also incorporated as baselines to be compared with our KGTrimmer method for recommendation.
  • Pop: We sort the entities according to the popularity, and we remove the entities with higher popularity. Here, the in-degree of entities is set as the measure of popularity.
  • KGRS-IMP: We adopt the GNN framework in this work, and we set $\boldsymbol{s}=\mathbf{1}$ to neglect the importance scores. As a result, we prune the edges based on lower trained embedding norm values.
  • RotatE(Sun et al., 2019): It is a link prediction method in knowledge graph embedding that defines each relation as a rotation from the head entity to the tail entity in the complex vector space. Here, the edges with lower prediction scores will be removed in our task.
  • KGAT(Wang et al., 2019a): This work explicitly models the high-order connectivities in KG and employs an attention mechanism to tell the importance of neighbors. To adapt our task, we directly sort the learned edge scores and prune the lower ones, resulting in a pruned KG.
  • KGist(Belth et al., 2020): The state-of-the-art KG summarization method innovatively adopts the Minimum Description Length principle (MDL) to inductively compress the original KG.
  • CAGED(Zhang et al., 2022): It is a state-of-the-art error detection method in KG that predicts the scores of triplets by multi-view contrastive learning. And we finally remove the anomaly entities with connected edges to get the pruned KG.
4.1.5. Reproductive Settings.
We implement our KGTrimmer in PyTorch. For a fair comparison, we fix the embedding size to 64 for all methods. We optimize the model parameters with Adam optimizer. Learning rate, regularization coefficient and dropout rate adhere to the settings of previous works(Wang et al., 2021a). In detail, the learning rate is searched amongst $\{1e^{-5},1e^{-4},1e^{-3}\}$, the regularization coefficient is tuned in $\{1e^{-6},1e^{-5},1e^{-4},1e^{-3}\}$. As for the GNN structure, the depth of GNN layers is selected from $\{1,2,3\}$, We leverage dropout operations to avoid over-fitting, where the node dropout and message dropout ratios are both tuned amongst $\{0.1,0.2,...,0.7\}$ separately. Furthermore, we choose at most $k\in\{20,50,100,200\}$ users for each entity to construct user-entity matrix. As for the transform function $\mathcal{T}$, we select from the identity function, average function and multiplicative function. Moreover, we adopt the early stopping strategy by performing premature stopping if the performance does not increase for 100 successive epochs.
4.2. Experimental Results
4.2.1. Overall Performance
In order to demonstrate the superiority of our proposed KGTrimmer methods for knowledge graph recommendation, we compare it with other prune baselines under two state-of-the-art knowledge graph recommendation backbones, and the results are shown in Table 2. From the results, we can get several observations.
(1) The performance of our proposed KGTrimmer surpasses most of baselines with the same pruning ratio under all datasets, demonstrating the effectiveness in preserving the crucial information of the knowledge graph for recommendation task. It is because collaborative filtering signals enable our model to better capture the influence of entities on users, while other baselines do not consider the role of users in the pruning task. And thanks to the dual-view module, our model can retain the filtered collective information and beneficial holistic information. (2) The advantages of our method become more evident as the pruning ratio increases. At a pruning ratio of first-class (90% in Amazon-book, 30% in Last-FM and Alibaba-Fashion), the performance of a few baselines remains comparable with our method, suggesting that they indeed filter some valueless or noisy entities. However, when pruning more triplets, our method significantly outperforms them because pruning solely relies on the KG itself has reached a performance bottleneck, while the advantages of our model in capturing collaborative filtering signals gradually become prominent. (3) Different backbones exhibit varying performance towards the quality of the KG. Specifically, KGIN can achieve better performance compared to KGRec at the same pruning ratio. We attribute this gap to the mask mechanism in KGRec which heightens the reliance on the quality of KG. (4) CAGED and KGist are found to be suboptimal in this task, especially in Amazon-book dataset. Although they perform well in noisy and error detection task, they are not originally designed for identifying valueless node and thus fail to recognize the true but redundant information.
Table 3. Time cost and performance comparison between the original KG and the pruned KG
Dataset Ratio Epoch Total Time R@20 N@20
Amazon-book 0% 579 5.5 days 0.1688 0.0915
90% 499 0.45 days 0.1667 0.0898
Last-FM 0% 509 3.15 days 0.0978 0.0848
30% 489 2.31 days 0.0966 0.0838
Alibaba-Fashion 0% 209 1.06 days 0.1146 0.0716
30% 219 0.85 days 0.1179 0.0738
4.2.2. Comparison Between Entire KG and the Pruned KG
The ultimate goal of pruning is to maintain the prediction accuracy while speeding up our prediction, especially in large-scale knowledge graph recommendation tasks. As a result, we compare the pruned knowledge graph generated from KGTrimmer with original one under the KGIN backbone. From the Table 3, we observe that (1) The performance trained on the pruned KG remains comparable with the whole KG, or even surpasses that on Alibaba-Fashion dataset. This observation verifies our motivation that the presence of numerous valueless nodes and noisy nodes in the KG may harm the recommendation performance. (2) KG pruning can lead to significant time efficiency improvements. It is worth noting that the training efficiency has increased by 12.2 times without any compromise in performance in Amazon-book dataset. Even though the number of user-item interactions is several times higher than the number of triplets in the Last-FM and Alibaba-Fashion dataset, KG pruning still offers substantial time efficiency gains. This phenomenon can be attributed to sophisticated structures designed on the KG in the existing models, which can boost performance but also increase the training time. (3) Different datasets contain varying number of valueless nodes and noisy nodes. To be specific, in the Amazon-book dataset, performance deterioration only occurs after pruning up to 95%, while in Last-FM, performance decline is seen at 30%, and in the Alibaba-Fashion dataset, a relatively good performance is still maintained even after 70% pruning.
4.2.3. Effect of Pruning Threshold
During experimental phase, we employ pruning ratio to ensure a fair comparison among baselines. In reality, it is rational to utilize threshold $\tau$ to prune the KG, namely all triplets below $\tau$ will be removed.
Figure 3. The final importance scores’ distribution in three datasets. The blue line and red line represent the distribution of entity scores and triplet scores, respectively.
To more intuitively display the selection of $\tau$, we have visualized the distribution graph of the final importance score. It should be highlighted that triplet scores are derived from the aggregation of entity scores referring to Equation 7, such as $\hat{s}(h,r,t)=\hat{s_{t}}$ , thereby implying that the observed discrepancies in their respective distributions are associated with the degree distribution of the entities. From Figure 3, we can draw the following observations: (1) All datasets contain vast number of triplets with low scores (less than 0.3), suggesting that non-informative information is a common issue across these datasets. (2)Despite the Amazon-book dataset presenting a limited number of entities with low scores, an unexpectedly large volume of triples have low scores. Examination of the graph structure revealed that seven entities with extremely low score contribute to half of all the triplets in the dataset. This finding confirms that a minority of valueless nodes consume the majority of computational resources.
In summary, selecting an appropriate $\tau$ can be directly applied to these three datasets. If one wish to retain as much useful information as possible, it is feasible to adopt a low $\tau$. Conversely, if one desire faster training efficiency, a larger $\tau$ is a better choice.
4.2.4. Ablation Study
In this part, we conduct the ablation study to show how the proposed dual-view importance evaluator affects the final performance of KGTrimmer in Table 4. Specifically, removing the collective view results in the dramatic decline of overall performance, undoubtedly proving the indispensable function of the collective view. On the other hand, the variant without holistic view also exhibits inferior performance , particularly in Amazon-Book dataset, which highlights the importance of capturing the impact of entities on recommendation tasks.
Table 4. Ablation studies that demonstrate the effectiveness of the model designs in Dual-View Importance Evaluator.
Amazon-book Last-FM Alibaba-Fashion
Recall NDCG Recall NDCG Recall NDCG
w/o col 0.1501 0.0790 0.0824 0.0714 0.1013 0.0611
w/o hol 0.1466 0.0778 0.0888 0.0788 0.1083 0.0668
all 0.1615 0.0861 0.0889 0.0789 0.1112 0.0694
Figure 4. Parameters sensitivity to show the how the $\gamma$ affect the performance.
4.2.5. Parameters Sensitivity
Moreover, we investigate the sensitivity of hyperparameters in this section.
The parameter $\gamma$ is primarily used to balance the weights between the two views. Specifically, when $\gamma=1$, only the impact of the collective view is considered, whereas when $\gamma=0$ , only the impact of the holistic view is taken into account. The results of these two aspects have already been demonstrated in the ablation studies. From Figure 4, we have arrived at the following conclusion : (1) Setting $\gamma=0.5$ services as a favorable trade-off for different datasets, indicating that both collective view and holistic view are beneficial to the task. 2) The Amazon-Book dataset experiences a dramatically decline at $\gamma$ values of 0.1 and 0.9, indicating that it relies heavily on both the collective and holistic views. Meanwhile, Alibaba-Fashion and Last-FM see a sharp drop at a $\gamma$ value of 0.1, demonstrating that the collective view is of crucial importance for them.
5. RELATED WORK
5.1. Knowledge-Aware Recommendation
Recent knowledge-aware recommendation literature can roughly fall into three categories: path-based, embedding-based and GNN-based models.
Path-based models (Catherine and Cohen, 2016; Hu et al., 2018; Ma et al., 2019; Sun et al., 2018; Wang et al., 2018, 2019b) focus on the patterns of the long-connectivity among users and entities via items. These patterns, such as meta-path (Hu et al., 2018; Jin et al., 2020), are delicately designed to capture the user preference on certain entities, and thus shed some light on the interpretation. However, these schemes are gradually fading out due to the heavy dependency on pattern quality.
Embedding-based models (Ai et al., 2018; Cao et al., 2019; Huang et al., 2018; Wang et al., 2020; Zhang et al., 2016) attempt to incorporate the semantic information in knowledge graph. They equip items with the KGE (knowledge graph embedding) techniques to discover latent relationships among entities. Despite achieving impressive performance on recommendation tasks, these models overlook the higher-order connectivity within the graph.
GNN-based models (Wang et al., 2021a, 2019c; Du et al., 2022; Wei et al., 2022) target at aggregating messages from neighbors to enrich the ego node embedding. which inherits the propagation mechanism of GNN(Zhou et al., 2020). Typically, KGIN (Wang et al., 2021a) proposes an end-to-end fashion to learn the latent user intents and provide more insight on interpretability. KGRec(Yang et al., 2023) proposes a rational-aware masking mechanism to identify the most important information in the KG, which mitigate the influence of the noisy and irrelevant triplets. Although they have achieved stunning success, all of them suffer from extremely high computational costs due to the cumbersome KG. Along this line, we propose a simple yet effective method to prune the undesirable nodes and maintain comparable performance on various models and datasets.
5.2. Knowledge Graph Pruning
Knowledge graph pruning techniques have garnered widespread attention due to the knowledge explosion issue in the information overload era(Tian et al., 2021). Nevertheless, extensive works confine themselves within the paradigm of subgraph extraction (Teru et al., 2020; Xu et al., 2020; Joshi and Urbani, 2020), which still retains a significant amount of irrelevant information, leading to wastage of information resources. Recently, Crumb-Trail (Faralli et al., 2018) tells the story in a different way, where it refines the taxonomic structure and permanently deletes the irrelevant edges. Notwithstanding, it only considers the structure of KGs, overlooking the role of collaborative signals from recommendation system.
Furthermore, researchers have devoted themselves to another line to slim the KGs, called Graph Summarization (GS), which compresses the vast KG into a compact one(Liu et al., 2018; Shabani et al., 2023; Doerr and Blenn, 2013; Sacenti et al., 2022; Belth et al., 2020; Kang et al., 2022; Brody et al., 2022; Goyal et al., 2020). One brunch of these approaches leverage machine learning or statistical rules to aggregate nodes. Some works (Doerr and Blenn, 2013; Sacenti et al., 2022) adopt clustering techniques to aggregate neighbors into a super node. Despite the prevalence of these methods, they still are incapable to incorporate the collaborative signals. KGist(Belth et al., 2020) tries to capture the meta-path rules according to the Minimum Description Length principle (MDL). By applying these rules, abnormal edges are identified and eliminated. Another stream of works turns to GNN for help(Kang et al., 2022; Brody et al., 2022; Goyal et al., 2020). For instance, PeGaSus (Kang et al., 2022) optimizes the summarization of large-scale graphs by considering both user preferences and interests, along with the structural properties of the graph. Regardless of remarkable achievements, KG summarization skills still face undesirable limitations. On the one hand, summarized nodes from the aggregated nodes may deviate from the facts that nodes in knowledge graph represent symbolic abstractions of entities from the real-world. On the other hand, they all put emphasis on the role of local neighbors, the signals from the holistic view are greatly neglected. Along this line, we launched an unexplored task knowledge graph pruning for the recommender system and proposed a novel dual-view-based framework to address the two issues above elegantly.
6. Conclusion
In this paper, we proposed a novel knowledge graph pruning model for recommendation which accelerates the training process and keep the performance stable. Specifically, we first designed a dual-view importance evaluator module to generate aggregated importance scores from collective and holistic views to assess the importance of each entity. Then, based on the importance scores, we constructed the importance-aware graph neural network to build the connection between recommendation system and evaluator module. Then, a new pruned knowledge graph was generated during the training process for downstream recommendation tasks with its lightweight and stable properties. Finally, extensive experiments on three real-world datasets demonstrated the effectiveness of the KGTrimmer compared with competitive baselines.
Acknowledgements.